OpenStudy (anonymous):

Find a one-parameter family of solutions of the differential equation dy/dt = 4y +1

OpenStudy (dumbcow):

separate variables \[\frac{dy}{4y+1} = dt\] integrate both sides \[\frac{1}{4} \ln (4y+1) = t +C\] \[4y+1 = e^{4t+C}\] \[y = \frac{ke^{4t} -1}{4}\]

OpenStudy (psymon):

I guess the whole "one-parameter" condition doesn't make it any different than a regular separation of variables and integrate O.o

OpenStudy (dumbcow):

yeah i think they are referring to the constant of integration...there is only 1 constant

OpenStudy (psymon):

Yeah, my best guess, too. Makes sense :P I haven't done too much with DEs, but theyre fun :D

OpenStudy (anonymous):

Cheers! My biggest problem always seems to be interpreting the questions.